# Joint Cooperative and Non-Cooperative Localization in WSNs with Distributed Scaled Proximal ADMM Algorithms

### v1

@@ Appendix A Solution to Remark 1
$=$ $\arg\min_{\begin{subarray}{c}\mathbf{p}_{i,j}^{+}=\mathbf{p}_{i,j}^{-}\\ \mathbf{q}_{i,j}^{+}=\mathbf{q}_{i,j}^{-}\\ \mathbf{x}_{i}=\mathbf{a}_{i},i\in\mathcal{A}\end{subarray}}\sum_{i\in\mathcal{N}}\bigg(\sum_{j\in\mathcal{N}_{i}}\bigg(\frac{1}{2}\|\mathbf{x}_{i}-\mathbf{p}_{i,j}^{+}\|^{2}-d_{i,j}\mathbf{v}_{i,j}^{T}(\mathbf{x}_{i}-\mathbf{p}_{i,j}^{+})$
$\quad+{\mathbf{\lambda}_{i,j}^{1}}^{T}(\mathbf{x}_{i}-\mathbf{p}_{i,j}^{-})+{\mathbf{\lambda}_{i,j}^{2}}^{T}(\mathbf{y}_{i}-\mathbf{q}_{i,j}^{-})+{\mathbf{\lambda}_{i,j}^{3}}^{T}(\mathbf{y}_{i}-\mathbf{q}_{i,j}^{+})\bigg)$
$\quad+\frac{1}{2}\|\mathbf{x}_{i}-\mathbf{y}_{i}\|^{2}-r_{i}\mathbf{u}_{i}^{T}(\mathbf{x}_{i}-\mathbf{y}_{i})\bigg)$
$=$ $\arg\min_{\begin{subarray}{c}\mathbf{p}_{i,j}^{+}=\mathbf{p}_{i,j}^{-}\\ \mathbf{q}_{i,j}^{+}=\mathbf{q}_{i,j}^{-}\\ \mathbf{x}_{i}=\mathbf{a}_{i},i\in\mathcal{A}\end{subarray}}\sum_{i\in(\mathcal{N}/\mathcal{A})}\bigg(\sum_{j\in\mathcal{N}_{i}}\bigg(\frac{1}{2}\|\mathbf{x}_{i}-\mathbf{p}_{j,i}^{-}\|^{2}-d_{i,j}\mathbf{v}_{i,j}^{T}(\mathbf{x}_{i}-\mathbf{p}_{j,i}^{-})$
$\quad+{\mathbf{\lambda}_{i,j}^{1}}^{T}(\mathbf{x}_{i}-\mathbf{p}_{i,j}^{-})+{\mathbf{\lambda}_{i,j}^{2}}^{T}(\mathbf{y}_{i}-\mathbf{q}_{i,j}^{-})+{\mathbf{\lambda}_{i,j}^{3}}^{T}(\mathbf{y}_{i}-\mathbf{q}_{j,i}^{-})\bigg)$
$\quad+\frac{1}{2}\|\mathbf{x}_{i}-\mathbf{y}_{i}\|^{2}-r_{i}\mathbf{u}_{i}^{T}(\mathbf{x}_{i}-\mathbf{y}_{i})\bigg)$
@@
$\quad+{\mathbf{\lambda}_{i,j}^{1}}^{T}(\mathbf{a}_{i}-\mathbf{p}_{i,j}^{-})+{\mathbf{\lambda}_{i,j}^{2}}^{T}(\mathbf{y}_{i}-\mathbf{q}_{i,j}^{-})+{\mathbf{\lambda}_{i,j}^{3}}^{T}(\mathbf{y}_{i}-\mathbf{q}_{j,i}^{-})\bigg)$
$\quad+\frac{1}{2}\|\mathbf{a}_{i}-\mathbf{y}_{i}\|^{2}-r_{i}\mathbf{u}_{i}^{T}(\mathbf{a}_{i}-\mathbf{y}_{i})\bigg).$ (D.18)
The final equality above holds by substituting the constraints $\mathbf{p}_{i,j}^{+}=\mathbf{p}_{j,i}^{-},\mathbf{q}_{i,j}^{+}=\mathbf{q}_{j,i}^{-}$ for all $i\in\mathcal{N},j\in\mathcal{N}_{i}$ and $\mathbf{x}_{i}=\mathbf{a}_{i},\forall i\in\mathcal{A}$ into the objective function. Using the KKT condition (43a), we conclude that $\mathbf{z}_{i}^{t}$ is the optimal solution of problem (D). From the optimality conditions of problem (D) with respect to $\mathbf{x}_{i}$ and $\mathbf{p}_{i,j}^{-}$, we obtain:
$\mathbf{0}$ $=\sum_{j\in\mathcal{N}_{i}}\big(\mathbf{x}_{i}^{t}-(\mathbf{p}_{j,i}^{-})^{t}-d_{i,j}\mathbf{v}_{i,j}^{t}+{\mathbf{\lambda}_{i,j}^{1}}^{t}\big)+\mathbf{x}_{i}^{t}-\mathbf{y}_{i}^{t}-r_{i}\mathbf{u}_{i}^{t},\quad\forall i\in(\mathcal{N}/\mathcal{A}),$
$\mathbf{0}$ $=\sum_{j\in\mathcal{N}_{i}}\big((\mathbf{p}_{i,j}^{-})^{t}-\mathbf{x}_{j}^{t}+d_{j,i}\mathbf{v}_{j,i}^{t}-{\mathbf{\lambda}_{i,j}^{1}}^{t}\big),\quad\forall i\in(\mathcal{N}/\mathcal{A})$
@@
$\mathbf{0}=\sum_{j\in\mathcal{N}_{i}}\left(\mathbf{x}_{i}^{t}-(\mathbf{z}_{j,i}^{-})^{t}-d_{i,j}\mathbf{v}_{i,j}^{t}+(\mathbf{z}_{i,j}^{-})^{t}-\mathbf{x}_{j}^{t}+d_{j,i}\mathbf{v}_{j,i}^{t}\right)+\mathbf{x}_{i}^{t}-\mathbf{y}_{i}^{t}-r_{i}\mathbf{u}_{i}^{t},\quad\forall i\in(\mathcal{N}/\mathcal{A}).$ (D.19)
@@
$\mathbf{0}=\sum_{j\in\mathcal{N}_{i}}\left(\mathbf{x}_{i}^{t}-\mathbf{x}_{j}^{t}-d_{i,j}\mathbf{v}_{i,j}^{t}+\mathbf{x}_{i}^{t}-\mathbf{x}_{j}^{t}+d_{j,i}\mathbf{v}_{j,i}^{t}\right)+\mathbf{x}_{i}^{t}-\mathbf{y}_{i}^{t}-r_{i}\mathbf{u}_{i}^{t},\quad\forall i\in(\mathcal{N}/\mathcal{A}).$ (D.20)
This indicates that $(\mathbf{x}^{t},\mathbf{y}_{1}^{t})$ satisfies equation (D.11).
Similarly, using the optimality conditions of problem (D) with respect to variables $\mathbf{y}_{i}$ and $\mathbf{q}_{i,j}^{-}$, we obtain:
$\begin{cases}\mathbf{0}&=\sum_{j\in\mathcal{N}_{i}}\big({\mathbf{\lambda}_{i,j}^{2}}^{t}+{\mathbf{\lambda}_{i,j}^{3}}^{t}\big)+\mathbf{y}_{i}^{t}-\mathbf{x}_{i}^{t}+r_{i}\mathbf{u}_{i}^{t},\quad\forall i\in(\mathcal{N}/\mathcal{A}),\\ \mathbf{0}&=\sum_{j\in\mathcal{N}_{i}}\big(-{\mathbf{\lambda}_{i,j}^{2}}^{t}-{\mathbf{\lambda}_{j,i}^{3}}^{t}\big),\quad\forall i\in(\mathcal{N}/\mathcal{A}),\end{cases}$
$\begin{cases}\mathbf{0}&=\sum_{j\in\mathcal{N}_{i}}\big({\mathbf{\lambda}_{i,j}^{2}}^{t}+{\mathbf{\lambda}_{i,j}^{3}}^{t}\big)+\mathbf{y}_{i}^{t}-\mathbf{a}_{i}^{t}+r_{i}\mathbf{u}_{i}^{t},\quad\forall i\in\mathcal{A},\\ \mathbf{0}&=\sum_{j\in\mathcal{N}_{i}}\big({-\mathbf{\lambda}_{i,j}^{2}}^{t}-{\mathbf{\lambda}_{j,i}^{3}}^{t}\big),\quad\forall i\in\mathcal{A}.\end{cases}$
@@
$\mathbf{0}$ $=\sum_{j\in\mathcal{N}_{i}}{\color[rgb]{0,0,0}\left({\mathbf{\lambda}_{i,j}^{3}}^{t}-{\mathbf{\lambda}_{j,i}^{3}}^{t}\right)}+\mathbf{y}_{i}^{t}-\mathbf{x}_{i}^{t}+r_{i}\mathbf{u}_{i}^{t},\quad\forall i\in(\mathcal{N}/\mathcal{A}),$
$\mathbf{0}$ $=\sum_{j\in\mathcal{N}_{i}}{\color[rgb]{0,0,0}\left({\mathbf{\lambda}_{i,j}^{3}}^{t}-{\mathbf{\lambda}_{j,i}^{3}}^{t}\right)}+\mathbf{y}_{i}^{t}-\mathbf{a}_{i}^{t}+r_{i}\mathbf{u}_{i}^{t},\quad\forall i\in\mathcal{A},$ (D.21)
@@
$\mathbf{0}=\sum_{i\in\mathcal{N}}\sum_{j\in\mathcal{N}_{i}}\left(\mathbf{\lambda}_{i,j}^{3,t}-\mathbf{\lambda}_{j,i}^{3,t}\right)+\sum_{i\in\mathcal{N}/\mathcal{A}}(\mathbf{y}_{i}^{t}-\mathbf{x}_{i}^{t}+r_{i}\mathbf{u}_{i}^{t})+\sum_{i\in\mathcal{A}}(\mathbf{y}_{i}^{t}-\mathbf{a}_{i}^{t}+r_{i}\mathbf{u}_{i}^{t}),$
@@
$\mathbf{0}=\sum_{i\in\mathcal{N}}\sum_{j\in\mathcal{N}_{i}}\left(\mathbf{\lambda}_{i,j}^{3,t}-\mathbf{\lambda}_{j,i}^{3,t}\right)$
Using the KKT condition (43a), the constraint $\mathbf{y}_{i}=\mathbf{y}_{j},\forall(i,j)\in\mathcal{E}$ holds. Given the sensor network is connected, it follows that all nodes reach consensus on the target position estimate, i.e., $\mathbf{y}_{i}^{t}=\mathbf{y}_{1}^{t}$ for all $i\in\mathcal{N}$. Thus:
$\mathbf{0}=\sum_{i\in\mathcal{N}/\mathcal{A}}(\mathbf{y}_{1}^{t}-\mathbf{x}_{i}^{t}+r_{i}\mathbf{u}_{i}^{t})+\sum_{i\in\mathcal{A}}(\mathbf{y}_{1}^{t}-\mathbf{a}_{i}^{t}+r_{i}\mathbf{u}_{i}^{t}),$
This indicates that $(\mathbf{x}^{t},\mathbf{y}_{1}^{t})$ satisfies equation (D.12). Hence, we obtain that $(\mathbf{x}^{t},\mathbf{y}_{1}^{t},\mathbf{w}^{t})$ is a critical point of problem (10). Consequently, it follows naturally that $(\mathbf{x}^{t},\mathbf{y}_{1}^{t})$ is a critical point of the original problem (2).
Appendix E (Proof of Lemma 3)
@@
$K:=(\mathcal{P}^{0}-\underline{\mathcal{P}})\cdot\max\big\{$ $\frac{1}{2},\frac{\kappa_{1}-1}{2},\frac{c(\kappa_{1}-1)}{2},\frac{\rho}{4},\frac{c(\kappa_{1}-6(2N_{\max}+1))}{2},\frac{c\kappa_{1}-6(N_{\max}+2)}{2c},$
$\frac{c\kappa_{2}\cdot\tilde{\tau}_{\min}}{2N_{\text{sum}}n(c+1)^{2}}-\frac{(2N_{\max}+1)c\kappa_{1}}{2},\frac{\rho}{4}-d^{2}_{\max}(\kappa_{1}+\kappa_{2})\big\}.$
@@
$\leq$ $d_{\max}\sqrt{N_{i}+2}\|\mathbf{w}_{i}^{s+1}-\mathbf{w}_{i}^{s}\|+\sqrt{2N_{i}+1}\|\bm{\lambda}_{i}^{s+1}-\bm{\lambda}_{i}^{s}\|$
$+(1+2c)(2+N_{\max})\|\mathbf{z}_{i}^{s+1}-\mathbf{z}_{i}^{s}\|,$ (E.2)
The last inequality holds because $\|\mathbf{H}_{i}^{T}\mathbf{H}_{i}\|=N_{i}+2$ and $\|\mathbf{A}_{i}^{T}\mathbf{A}_{i}\|=2N_{i}+1$, and based on (B.12).
Substituting equations (E) and (B) into equation (E) yields
$\|\bm{\eta}_{1}\|\leq\sqrt{\frac{K}{t-1}}\left(d_{\max}\sqrt{N_{i}+2}+\sqrt{(2N_{i}+1)[3(N_{i}+2)+3c^{2}(4N_{i}+2)+3c(1+2c)(2+N_{\max})]}\right)\leq\sqrt{\frac{M_{1}}{t-1}},$
@@
$M_{1}:=K\left(d_{\max}\sqrt{N_{i}+2}+\sqrt{(2N_{i}+1)[3(N_{i}+2)+3c^{2}(4N_{i}+2)+3c(1+2c)(2+N_{\max})]}\right)>0.$
@@
$\mathbf{w}_{i}^{s+1}=\arg\min G_{i}(\mathbf{z}_{i}^{s+1},\mathbf{w}_{i})+\frac{\rho}{2}\|\mathbf{w}_{i}-\mathbf{w}_{i}^{s}\|^{2}+\delta_{\mathbf{B}_{N_{i}+1}}(\mathbf{w}_{i}).$
@@
$\mathbf{0}\in\nabla_{\mathbf{w}_{i}}G_{i}(\mathbf{z}_{i}^{s+1},\mathbf{w}_{i}^{s+1})+\rho(\mathbf{w}_{i}^{s+1}-\mathbf{w}_{i}^{s})+\partial\delta_{\mathbf{B}_{N_{i}+1}}(\mathbf{w}_{i}^{s+1}).$
@@
$\bm{\eta}_{2}\in\nabla_{\mathbf{w}_{i}}G_{i}(\mathbf{z}_{i}^{s+1},\mathbf{w}_{i}^{s+1})+\partial\delta_{\mathbf{B}_{N_{i}+1}}(\mathbf{w}_{i}^{s+1}).$
Combining this with inequality (E), we get
$\|\bm{\eta}_{2}\|\leq\sqrt{\frac{K}{t-1}}\cdot\rho=\frac{M_{2}}{\sqrt{t}},$
@@
$M_{2}:=K\rho^{2}>0.$
Therefore, Lemma 3 holds for $\bm{\eta}_{1},\bm{\eta}_{2}$, and $M:=\max\{M_{1},M_{2}\}$, which implies that $(\mathbf{z}^{t},\mathbf{w}^{t},\bm{\lambda}^{t})$ is an $M/\sqrt{t-1}$-solution to problem (24). This shows that the proposed algorithm converges to a Karush-Kuhn-Tucker point of problem (24) at a sublinear rate.
Appendix F (Proof of Theorem 2)

### v2

@@ A.1 Solution to Remark 2
$=$ $\arg\min_{\begin{subarray}{c}\mathbf{p}_{i,j}^{+}=\mathbf{p}_{i,j}^{-}\\ \mathbf{q}_{i,j,{\color[rgb]{0,0,0}k}}^{+}=\mathbf{q}_{i,j,{\color[rgb]{0,0,0}k}}^{-},\forall k\in\mathcal{M}\\ \mathbf{x}_{i}=\mathbf{a}_{i},\forall i\in\mathcal{A}\end{subarray}}\sum_{i\in\mathcal{N}}\bigg(\sum_{j\in\mathcal{N}_{i}}\bigg(\frac{1}{2}\|\mathbf{x}_{i}-\mathbf{p}_{i,j}^{+}\|^{2}-d_{i,j}\mathbf{v}_{i,j}^{T}(\mathbf{x}_{i}-\mathbf{p}_{i,j}^{+})$
$\quad+{\mathbf{\lambda}_{i,j}^{0}}^{T}(\mathbf{x}_{i}-\mathbf{p}_{i,j}^{-})+\sum_{k\in\mathcal{M}}\big({\mathbf{\lambda}_{i,j}^{2k-1}}^{T}(\mathbf{y}_{i,{\color[rgb]{0,0,0}k}}-\mathbf{q}_{i,j,{\color[rgb]{0,0,0}k}}^{-})+{\mathbf{\lambda}_{i,j}^{2k}}^{T}(\mathbf{y}_{i,{\color[rgb]{0,0,0}k}}-\mathbf{q}_{i,j,{\color[rgb]{0,0,0}k}}^{+})\big)\bigg)$
$\quad+\sum_{k\in\mathcal{M}_{i}}\big(\frac{1}{2}\|\mathbf{x}_{i}-\mathbf{y}_{i,{\color[rgb]{0,0,0}k}}\|^{2}-r_{i,{\color[rgb]{0,0,0}k}}\mathbf{u}_{i,{\color[rgb]{0,0,0}k}}^{T}(\mathbf{x}_{i}-\mathbf{y}_{i,{\color[rgb]{0,0,0}k}})\big)\bigg)$
$=$ $\arg\min_{\begin{subarray}{c}\mathbf{p}_{i,j}^{+}=\mathbf{p}_{i,j}^{-}\\ \mathbf{q}_{i,j,{\color[rgb]{0,0,0}k}}^{+}=\mathbf{q}_{i,j,{\color[rgb]{0,0,0}k}}^{-}\\ \mathbf{x}_{i}=\mathbf{a}_{i},i\in\mathcal{A}\end{subarray}}\sum_{i\in(\mathcal{N}/\mathcal{A})}\bigg(\sum_{j\in\mathcal{N}_{i}}\bigg(\frac{1}{2}\|\mathbf{x}_{i}-\mathbf{p}_{j,i}^{-}\|^{2}-d_{i,j}\mathbf{v}_{i,j}^{T}(\mathbf{x}_{i}-\mathbf{p}_{j,i}^{-})$
$\quad+{\mathbf{\lambda}_{i,j}^{0}}^{T}(\mathbf{x}_{i}-\mathbf{p}_{i,j}^{-})+\sum_{k\in\mathcal{M}}\big({\mathbf{\lambda}_{i,j}^{2k-1}}^{T}(\mathbf{y}_{i,{\color[rgb]{0,0,0}k}}-\mathbf{q}_{i,j,{\color[rgb]{0,0,0}k}}^{-})+{\mathbf{\lambda}_{i,j}^{2k}}^{T}(\mathbf{y}_{i,{\color[rgb]{0,0,0}k}}-\mathbf{q}_{j,i,{\color[rgb]{0,0,0}k}}^{-})\big)\bigg)$
$\quad+\sum_{k\in\mathcal{M}_{i}}\big(\frac{1}{2}\|\mathbf{x}_{i}-\mathbf{y}_{i,{\color[rgb]{0,0,0}k}}\|^{2}-r_{i,{\color[rgb]{0,0,0}k}}\mathbf{u}_{i,{\color[rgb]{0,0,0}k}}^{T}(\mathbf{x}_{i}-\mathbf{y}_{i,{\color[rgb]{0,0,0}k}})\big)\bigg)$
@@
$\quad+{\mathbf{\lambda}_{i,j}^{0}}^{T}(\mathbf{a}_{i}-\mathbf{p}_{i,j}^{-})+\sum_{k\in\mathcal{M}}\big({\mathbf{\lambda}_{i,j}^{2k-1}}^{T}(\mathbf{y}_{i,{\color[rgb]{0,0,0}k}}-\mathbf{q}_{i,j,{\color[rgb]{0,0,0}k}}^{-})+{\mathbf{\lambda}_{i,j}^{2k}}^{T}(\mathbf{y}_{i,{\color[rgb]{0,0,0}k}}-\mathbf{q}_{j,i,{\color[rgb]{0,0,0}k}}^{-})\big)\bigg)$
$\quad+\sum_{k\in\mathcal{M}_{i}}\big(\frac{1}{2}\|\mathbf{a}_{i}-\mathbf{y}_{i,{\color[rgb]{0,0,0}k}}\|^{2}-r_{i,{\color[rgb]{0,0,0}k}}\mathbf{u}_{i,{\color[rgb]{0,0,0}k}}^{T}(\mathbf{a}_{i}-\mathbf{y}_{i,{\color[rgb]{0,0,0}k}})\big)\bigg).$ (D.18)
The final equality above holds by substituting the constraints $\mathbf{p}_{i,j}^{+}=\mathbf{p}_{j,i}^{-},\mathbf{q}_{i,j,{\color[rgb]{0,0,0}k}}^{+}=\mathbf{q}_{j,i,{\color[rgb]{0,0,0}k}}^{-}$ for all $i\in\mathcal{N},j\in\mathcal{N}_{i}$ and $\mathbf{x}_{i}=\mathbf{a}_{i},\forall i\in\mathcal{A}$ into the objective function. Using the KKT condition (42a), we conclude that $\mathbf{z}_{i}^{t}$ is the optimal solution of problem (A.4). From the optimality conditions of problem (A.4) with respect to $\mathbf{x}_{i}$ and $\mathbf{p}_{i,j}^{-}$, we obtain:
$\mathbf{0}$ $=\sum_{j\in\mathcal{N}_{i}}\big(\mathbf{x}_{i}^{t}-(\mathbf{p}_{j,i}^{-})^{t}-d_{i,j}\mathbf{v}_{i,j}^{t}+{\mathbf{\lambda}_{i,j}^{0}}^{t}\big)+\sum_{k\in\mathcal{M}_{i}}\big(\mathbf{x}_{i}^{t}-\mathbf{y}_{i,{\color[rgb]{0,0,0}k}}^{t}-r_{i,{\color[rgb]{0,0,0}k}}\mathbf{u}_{i,{\color[rgb]{0,0,0}k}}^{t}\big),\quad\forall i\in(\mathcal{N}/\mathcal{A}),$
$\mathbf{0}$ $=\sum_{j\in\mathcal{N}_{i}}\big((\mathbf{p}_{i,j}^{-})^{t}-\mathbf{x}_{j}^{t}+d_{j,i}\mathbf{v}_{j,i}^{t}-{\mathbf{\lambda}_{i,j}^{0}}^{t}\big),\quad\forall i\in(\mathcal{N}/\mathcal{A})$
@@
$\mathbf{0}=\sum_{j\in\mathcal{N}_{i}}\left(\mathbf{x}_{i}^{t}-(\mathbf{p}_{j,i}^{-})^{t}-d_{i,j}\mathbf{v}_{i,j}^{t}+(\mathbf{p}_{i,j}^{-})^{t}-\mathbf{x}_{j}^{t}+d_{j,i}\mathbf{v}_{j,i}^{t}\right)+\sum_{k\in\mathcal{M}_{i}}\big(\mathbf{x}_{i}^{t}-\mathbf{y}_{i,{\color[rgb]{0,0,0}k}}^{t}-r_{i,{\color[rgb]{0,0,0}k}}\mathbf{u}_{i,{\color[rgb]{0,0,0}k}}^{t}\big),\quad\forall i\in(\mathcal{N}/\mathcal{A}).$ (D.19)
@@
$\mathbf{0}=\sum_{j\in\mathcal{N}_{i}}\left(\mathbf{x}_{i}^{t}-\mathbf{x}_{j}^{t}-d_{i,j}\mathbf{v}_{i,j}^{t}+\mathbf{x}_{i}^{t}-\mathbf{x}_{j}^{t}+d_{j,i}\mathbf{v}_{j,i}^{t}\right)+\sum_{k\in\mathcal{M}_{i}}\big(\mathbf{x}_{i}^{t}-\mathbf{y}_{i,{\color[rgb]{0,0,0}k}}^{t}-r_{i,{\color[rgb]{0,0,0}k}}\mathbf{u}_{i,{\color[rgb]{0,0,0}k}}^{t}\big),\quad\forall i\in(\mathcal{N}/\mathcal{A}).$ (D.20)
This indicates that $(\mathbf{x}^{t},\mathbf{y}_{i}^{t})$ satisfies equation (D.11).
Similarly, using the optimality conditions of problem (A.4) with respect to variables $\mathbf{y}_{i,{\color[rgb]{0,0,0}k}}$ and $\mathbf{q}_{i,j,{\color[rgb]{0,0,0}k}}^{-}$, we obtain:
$\begin{cases}\mathbf{0}&=\sum_{j\in\mathcal{N}_{i}}\sum_{k\in\mathcal{M}}\big({\mathbf{\lambda}_{i,j}^{2k-1}}^{t}+{\mathbf{\lambda}_{i,j}^{2k}}^{t}\big)+\sum_{k\in\mathcal{M}_{i}}\big(\mathbf{y}_{i,{\color[rgb]{0,0,0}k}}^{t}-\mathbf{x}_{i}^{t}+r_{i,{\color[rgb]{0,0,0}k}}\mathbf{u}_{i,{\color[rgb]{0,0,0}k}}^{t}\big),\quad\forall i\in(\mathcal{N}/\mathcal{A}),\\ \mathbf{0}&=\sum_{j\in\mathcal{N}_{i}}\big(-{\mathbf{\lambda}_{i,j}^{2k-1}}^{t}-{\mathbf{\lambda}_{j,i}^{2k}}^{t}\big),\quad\forall i\in(\mathcal{N}/\mathcal{A}),\end{cases}$
$\begin{cases}\mathbf{0}&=\sum_{j\in\mathcal{N}_{i}}\sum_{k\in\mathcal{M}}\big({\mathbf{\lambda}_{i,j}^{2k-1}}^{t}+{\mathbf{\lambda}_{i,j}^{2k}}^{t}\big)+\sum_{k\in\mathcal{M}_{i}}\big(\mathbf{y}_{i,{\color[rgb]{0,0,0}k}}^{t}-\mathbf{a}_{i}^{t}+r_{i,{\color[rgb]{0,0,0}k}}\mathbf{u}_{i,{\color[rgb]{0,0,0}k}}^{t}\big),\quad\forall i\in\mathcal{A},\\ \mathbf{0}&=\sum_{j\in\mathcal{N}_{i}}\big({-\mathbf{\lambda}_{i,j}^{2k-1}}^{t}-{\mathbf{\lambda}_{j,i}^{2k}}^{t}\big),\quad\forall i\in\mathcal{A}.\end{cases}$
@@
$\mathbf{0}$ $=\sum_{j\in\mathcal{N}_{i}}\sum_{k\in\mathcal{M}}{\color[rgb]{0,0,0}\left({\mathbf{\lambda}_{i,j}^{2k}}^{T}-{\mathbf{\lambda}_{j,i}^{2k}}^{T}\right)}+\sum_{k\in\mathcal{M}_{i}}\big(\mathbf{y}_{i,{\color[rgb]{0,0,0}k}}^{t}-\mathbf{x}_{i}^{t}+r_{i,{\color[rgb]{0,0,0}k}}\mathbf{u}_{i}^{t}\big),\quad\forall i\in(\mathcal{N}/\mathcal{A}),$
$\mathbf{0}$ $=\sum_{j\in\mathcal{N}_{i}}\sum_{k\in\mathcal{M}}\big({\color[rgb]{0,0,0}\left({\mathbf{\lambda}_{i,j}^{2k}}^{T}-{\mathbf{\lambda}_{j,i}^{2k}}^{T}\right)}+\sum_{k\in\mathcal{M}_{i}}\big(\mathbf{y}_{i,{\color[rgb]{0,0,0}k}}^{t}-\mathbf{a}_{i}^{t}+r_{i,{\color[rgb]{0,0,0}k}}\mathbf{u}_{i,{\color[rgb]{0,0,0}k}}^{t}\big),\quad\forall i\in\mathcal{A},$ (D.21)
@@
$\mathbf{0}=\sum_{i\in\mathcal{N}}\sum_{j\in\mathcal{N}_{i}}\sum_{k\in\mathcal{M}}\left(\mathbf{\lambda}_{i,j}^{2k,t}-\mathbf{\lambda}_{j,i}^{2k,t}\right)+\sum_{i\in\mathcal{N}/\mathcal{A}}\sum_{k\in\mathcal{M}_{i}}(\mathbf{y}_{i,{\color[rgb]{0,0,0}k}}^{t}-\mathbf{x}_{i}^{t}+r_{i,{\color[rgb]{0,0,0}k}}\mathbf{u}_{i,k}^{t})+\sum_{i\in\mathcal{A}}\sum_{k\in\mathcal{M}_{i}}(\mathbf{y}_{i,{\color[rgb]{0,0,0}k}}^{t}-\mathbf{a}_{i}^{t}+r_{i,{\color[rgb]{0,0,0}k}}\mathbf{u}_{i,k}^{t}),$
@@
$\mathbf{0}=\sum_{i\in\mathcal{N}}\sum_{j\in\mathcal{N}_{i}}\sum_{k\in\mathcal{M}}\left(\mathbf{\lambda}_{i,j}^{2k,t}-\mathbf{\lambda}_{j,i}^{2k,t}\right)$
Using the KKT condition (42a), the constraint $\mathbf{y}_{i,{\color[rgb]{0,0,0}k}}=\mathbf{y}_{j,k},\forall(i,j)\in\mathcal{E}$ holds. Given the sensor network is connected, it follows that all nodes reach consensus on the target position estimate, i.e., $\mathbf{y}_{i,{\color[rgb]{0,0,0}k}}^{t}=\mathbf{y}_{1,k}^{t}$ for all $i\in\mathcal{N}$. Thus:
$\mathbf{0}=\sum_{i\in\mathcal{N}/\mathcal{A}}\sum_{k\in\mathcal{M}_{i}}(\mathbf{y}_{1,k}^{t}-\mathbf{x}_{i}^{t}+r_{i,{\color[rgb]{0,0,0}k}}\mathbf{u}_{i,k}^{t})+\sum_{i\in\mathcal{A}}\sum_{k\in\mathcal{M}_{i}}(\mathbf{y}_{1,k}^{t}-\mathbf{a}_{i}^{t}+r_{i,{\color[rgb]{0,0,0}k}}\mathbf{u}_{i,k}^{t}),$
This indicates that $(\mathbf{x}^{t},\mathbf{y}_{1}^{t})$ satisfies equation (D.12), where $\mathbf{y}_{1}^{t}=\text{vec}(\mathbf{y}_{1,k},k\in\mathcal{M})$. Hence, we obtain that $(\mathbf{x}^{t},\mathbf{y}_{1}^{t},\mathbf{w}^{t})$ is a critical point of problem (9). Consequently, it follows naturally that $(\mathbf{x}^{t},\mathbf{y}_{1}^{t})$ is a critical point of the original problem (2).
A.5 (Proof of Lemma 3)
@@
$K:=(\mathcal{P}^{0}-\underline{\mathcal{P}})\cdot\max\big\{$ $\frac{1}{2},\frac{\kappa_{1}-1}{2},\frac{c(\kappa_{1}-1)}{2},\frac{\rho}{4},\frac{c(\kappa_{1}-6(2N_{\max}+1))}{2},\frac{c\kappa_{1}-6(N_{\max}+M_{\max}+1)}{2c},$
$\frac{c\kappa_{2}\cdot\tilde{\tau}_{\min}}{2(2M+1)N_{\text{sum}}n(c+1)^{2}}-\frac{(2N_{\max}+1)c\kappa_{1}}{2},\frac{\rho}{4}-d^{2}_{\max}(\kappa_{1}+\kappa_{2})\big\}.$
@@
$\leq$ $d_{\max}\sqrt{N_{i}+M_{i}+1}\|\mathbf{w}_{i}^{s+1}-\mathbf{w}_{i}^{s}\|+\sqrt{2N_{i}+1}\|\bm{\lambda}_{i}^{s+1}-\bm{\lambda}_{i}^{s}\|$
$+(1+c)(2+N_{\max})\|\mathbf{z}_{i}^{s+1}-\mathbf{z}_{i}^{s}\|,$ (E.2)
The last inequality holds because $\|\mathbf{H}_{i}^{T}\mathbf{H}_{i}\|=N_{i}+M_{i}+1$ and $\|\mathbf{A}_{i}^{T}\mathbf{A}_{i}\|=2N_{i}+1$, and based on (B.12).
Substituting equations (A.5) and (A.2) into equation (A.5) yields
$\|\bm{\eta}_{1}\|$ $\leq\sqrt{\frac{K}{t-1}}\left(d_{\max}\sqrt{N_{i}+M_{i}+1}+\sqrt{(2N_{i}+1)[3(N_{i}+M_{i}+1)+3c^{2}(4N_{i}+2)+3c(1+c)(1+M_{\max}+N_{\max})]}\right)$
$\leq\sqrt{\frac{S_{1}}{t-1}},$
@@
$S_{1}:=K\left(d_{\max}\sqrt{N_{i}+M_{i}+1}+\sqrt{(2N_{i}+1)[3(N_{i}+M_{i}+1)+3c^{2}(4N_{i}+2)+3c(1+c)(1+M_{\max}+N_{\max})]}\right)>0.$
@@
$\mathbf{w}_{i}^{s+1}=\arg\min G_{i}(\mathbf{z}_{i}^{s+1},\mathbf{w}_{i})+\frac{\rho}{2}\|\mathbf{w}_{i}-\mathbf{w}_{i}^{s}\|^{2}+\delta_{\mathbf{B}^{N_{i}+M}}(\mathbf{w}_{i}).$
@@
$\mathbf{0}\in\nabla_{\mathbf{w}_{i}}G_{i}(\mathbf{z}_{i}^{s+1},\mathbf{w}_{i}^{s+1})+\rho(\mathbf{w}_{i}^{s+1}-\mathbf{w}_{i}^{s})+\partial\delta_{\mathbf{B}^{N_{i}+M}}(\mathbf{w}_{i}^{s+1}).$
@@
$\bm{\eta}_{2}\in\nabla_{\mathbf{w}_{i}}G_{i}(\mathbf{z}_{i}^{s+1},\mathbf{w}_{i}^{s+1})+\partial\delta_{\mathbf{B}^{N_{i}+M}}(\mathbf{w}_{i}^{s+1}).$
Combining this with inequality (A.5), we get
$\|\bm{\eta}_{2}\|\leq\sqrt{\frac{K}{t-1}}\cdot\rho=\frac{S_{2}}{\sqrt{t}},$
@@
$S_{2}:=K\rho^{2}>0.$
Therefore, Lemma 3 holds for $\bm{\eta}_{1},\bm{\eta}_{2}$, and $S:=\max\{S_{1},S_{2}\}$, which implies that $(\mathbf{z}^{t},\mathbf{w}^{t},\bm{\lambda}^{t})$ is an $S/\sqrt{t-1}$-solution to problem (23). This shows that the proposed algorithm converges to a Karush-Kuhn-Tucker point of problem (23) at a sublinear rate.
A.6 (Proof of Theorem 2)